# Eugene Lerman

### v1

@@ Appendix A Computing $\operatorname{Spec}(\mathscr{A})$
A.i. $\operatorname{Spec}(\mathscr{A})$ for a finitely generated $C^{\infty}$-ring $\mathscr{A}$
@@
Definition A.1.
A $C^{\infty}$-ring $\mathscr{A}$ is finitely generated if there is a surjective map $\varpi:C^{\infty}(\mathbb{R}^{N})\to\mathscr{A}$ of $C^{\infty}$-rings. (And then $\varpi(x_{1}),\ldots\varpi(x_{N})\in\mathscr{A}$ are the generators of $\mathscr{A}$.)
@@
A.ii. $\operatorname{Spec}(\mathscr{A})$ for a point determined $C^{\infty}$-ring $\mathscr{A}$
@@ A $C^{\infty}$-ring $\mathscr{A}$ is point determined if for every element $a\in\mathscr{A}$
$x(a)=0\textrm{ for all }x\in\operatorname{Hom}(\mathscr{A},\mathbb{R})\equiv X_{\mathscr{A}}\quad\Rightarrow a=0.$
@@
(A.2) $\mathscr{A}\to C^{0}(X_{\mathscr{A}}),\quad a\mapsto a_{*}$
@@
Suppose $a\in\bigcap_{x\in U}I_{x}$. Then for any $x\in U$, we have $a\in I_{x}$, which means that there is $d\in\mathscr{A}$ so that $x(d)\not=0$ and $ad=0$. Since $0\not=x(d)=d_{*}(x)$ and $d_{*}$ is continuous there is a neighborhood $V$ of $x$ so that $d_{*}(y)\not=0$ for all $y\in V$. Since $ad=0$, $0=(ad)_{*}=a_{*}d_{*}$. Hence $a_{*}|_{V}=0$. Since $x\in U$ is arbitrary, $a_{*}|_{U}=0$. Therefore
@@
$0=d_{*}a_{*}=(da)_{*}$
Since the map (A.2) is injective ($\mathscr{A}$ is point determned), it follows that $da=0$ Since $x(d)=x_{*}(d)\not=0$ it follows that $a\in I_{x}$. Since $x\in U$ is arbitrary, $a\in\bigcap_{x\in U}I_{x}$. Therefore $\mathfrak{s}_{a}|_{U}=0$ and $\psi_{U}$ is injective. ∎
@@
$I(Z_{J})=\{f\in C^{\infty}(\mathbb{R}^{n})\mid f(p)=0\textrm{ for all }p\in Z_{J}$
@@
(A.4) $\overline{\operatorname{\mathsf{ev}}}:Z_{J}\to\operatorname{Hom}(C^{\infty}(\mathbb{R}^{n}/J),\mathbb{R}),\quad\overline{\operatorname{\mathsf{ev}}}_{p}(f+J)=f(p).$
@@
for all $p\in Z_{J}$. Hence $x(f+J)=0$ for all $x\in\operatorname{Hom}(\mathscr{A},\mathbb{R})$. Since $\mathscr{A}$ is point determined $f+J=0$ in $\mathscr{A}=C^{\infty}(\mathbb{R}^{n})/J$. Thus $f\in J$. ∎
It follows from A.i that if $\mathscr{A}=C^{\infty}(\mathbb{R}^{n})/J$ is point determined then $X_{\mathscr{A}}=Z_{J}$ and $\mathscr{O}_{A}$ is given by
@@
A.iii. The spectrum of a $C^{\infty}$-ring $\operatorname{Spec}(C^{\infty}(M)^{G})$ of invariant functions.
@@
If $f\in C^{\infty}(M)^{G}$ and $x(f)=0$ for all $x\in\operatorname{Hom}(C^{\infty}(M)^{G},\mathbb{R})$, then $f(p)=0$ for all $p\in M$, and then $f=0$. ∎
@@
(A.5) $\overline{\operatorname{\mathsf{ev}}}:M/G\to X_{C^{\infty}(M)^{G}},\quad\overline{\operatorname{\mathsf{ev}}}(G\cdot p)=\operatorname{\mathsf{ev}}_{p}.$
In general there is no reason for the map $\overline{\operatorname{\mathsf{ev}}}$ to be a bijection. If invariant functions separate orbits (which happens when the action is proper) then $\overline{\operatorname{\mathsf{ev}}}$ is injective. In general it is not. For instance consider the action of the reals $\mathbb{R}$ on the torus $\mathbb{T}^{2}$ by irrational flow (cf. Subsection 4.i). Then the ring of invariants $C^{\infty}(\mathbb{T}^{2})^{\mathbb{R}}$ consists of constant functions. Consequently it has exactly one $\mathbb{R}$ point and the map $\overline{\operatorname{\mathsf{ev}}}:\mathbb{T}^{2}/\mathbb{R}\to X_{C^{\infty}(\mathbb{T}^{2})^{\mathbb{R}}}$ is not injective. We now argue that for proper actions the map $\overline{\operatorname{\mathsf{ev}}}$ is a homeomorphism. We do it in a series of lemmas.
Lemma A.7.
@@
Remark A.8.
Let $G\times M\to M$ be an action of a Lie group $G$ on a manifold $M$. Let $q:M\to M/G$ denote the orbit/quotient map. By the universal property of $q$ for any $G$-invariant function $f:M\to\mathbb{R}$ there exists a unique continuous function $\overline{f}:M/G\to\mathbb{R}$ with $f=\overline{f}\circ q$. We thus get an injective map
(A.6) $\Upsilon:C^{\infty}(M)^{G}\to C^{0}(M/G),\qquad f\mapsto\overline{f}$
@@
(A.7) $C^{\infty}(M/G):=\Upsilon(C^{\infty}(M)^{G})\qquad$
@@
Lemma A.9.
Let $G\times M\to M$ be a proper action of a Lie group $G$ on a manifold $M$. Then the $C^{\infty}$-subring $C^{\infty}(M/G)$ of $C^{0}(M/G)$ defined in Remark A.8 contains a proper function. Consequently the map
@@
In the course of the proof of Theorem 4.3.1 in [Pa] Palais constructs a sequence of invariant functions $\{f_{n}\}_{n\in\mathbb{N}}\subset C^{\infty}(M)^{G}$ with the property that their supports form a locally finite cover of $M$, that $f_{n}(M)\subset[0,1]$ for all $n$, $\sum f_{n}=1$ and that sets $q(\operatorname{supp}(f_{n}))\subset M/G$ are compact (here as before $q:M\to M/G$ is the quotient map). Let $\overline{f}_{n}$ denote the image of $f_{n}$ in $C^{0}(M/G)$ under the map $\Upsilon:C^{\infty}(M)^{G}\to C^{0}(M/G)$ (cf. Remark A.8). Then the function
$h=\sum_{n=1}^{\infty}n\overline{f}_{n})$
is the desired proper function (cf. proof of Proposition 2.28 in [Lee]). ∎
Lemma A.10.
@@
As we observed in Remark A.8 the map $\Upsilon:C^{\infty}(M)^{G}\to C^{\infty}(M/G)$ is an isomorphism of $C^{\infty}$-rings, hence $X_{C^{\infty}(M)^{G}}$ and $X_{C^{\infty}(M/G}$ are homeomorphic. By Lemma A.9 the evaluation map
@@
is a bijection. We need to show that $\operatorname{\mathsf{ev}}$ is a homeomorphism. By Lemma A.11 below the topology on $M/G$ is generated by the sets of the form $\{h\not=0\}:=\{p\in M/G\mid h(p)\not=0\}$, $h\in C^{\infty}(M/G)$. On the other hand the topology on $X_{C^{\infty}(M/G)}$ is generated by the sets of the form
@@
Lemma A.11.
Let $G\times M\to M$ be a proper action of a Lie group $G$ on a manifold $M$, $q:M\to M/G$ the quotient map and $C^{\infty}(M/G)\subset C^{0}(M/G)$ defined by (A.7). The quotient topology on $M/G$ is the smallest topology making all the functions in $C^{\infty}(M/G)$ continuous. That is, the quotient topology is generated by the sets of the form $\{h\not=0\}$ for all $h\in C^{\infty}(M/G)$.
@@
[…]
Lemma A.12.
@@
By Lemma A.10 the space of points $X_{C^{\infty}(M)^{G}}$ is homeomorphic to the orbit space $M/G$ by way of the map
@@
(A.8) $(\overline{\operatorname{\mathsf{ev}}})^{*}:C^{0}(X_{C^{\infty}(M)^{G}})\to C^{0}(M/G),\quad h\mapsto h\circ\overline{\operatorname{\mathsf{ev}}}$
@@
By Lemma A.6 the $C^{\infty}$-ring $C^{\infty}(M)^{G}$ is point determined, hence can be identified with a $C^{\infty}$-subring of $C^{0}(X_{C^{\infty}(M)^{G}})$ by way of
(A.9) $(\,\,)_{*}:C^{\infty}(M)^{G}\to C^{0}(X_{C^{\infty}(M)^{G}}),\quad f\mapsto f_{*}$
@@
the composite of (A.9) and (A.8) is the map $\Upsilon:C^{\infty}(M)^{G}\to C^{0}(M/G)$ that sends the invariant function $f\in C^{\infty}(M)^{G}$ to the unique function $\overline{f}:M/G\to\mathbb{R}$ with $\overline{f}\circ q=f$ (cf. Lemma A.8). Lemma A.3 now implies the desired result. ∎
A.iv. $\operatorname{Spec}(C^{\infty}(M)^{G})$ when the action $G\times M\to M$ is proper and $C^{\infty}(M)^{G}$ is finitely generated.
We now put together Subsections A.i and A.iii. Suppose now $G\times M\to M$ is a proper action and the $C^{\infty}$-ring of invariants $C^{\infty}(M)^{G}$ is finitely generated (hence there is a surjective map $\varpi:C^{\infty}(\mathbb{R}^{N})\to C^{\infty}(M)^{G}$ for some $N$). Then, one one hand,
@@
Appendix B A remark on the Cushman-Weinstein conjecture
Suppose $M$ is a symplectic manifold with a proper Hamiltonian action of a Lie group $G$ and an associated equivariant moment map $\mu:M\to\mathfrak{g}^{\vee}$. Suppose further that the $C^{\infty}$-ring $C^{\infty}(M)^{G}$ is finitely generated. Then the $C^{\infty}$-ring of invariants $C^{\infty}(M)^{G}$ is Poisson and consequently $\operatorname{Spec}(C^{\infty}(M)^{G})\simeq(M/G,C^{\infty}_{M/G})$ is a Poisson $C^{\infty}$-scheme. Moreover as we saw in A.iv a choice of a surjective map $\varpi:C^{\infty}(\mathbb{R}^{N})\to C^{\infty}(M)^{G}$ (i.e., a choice of generators of $C^{\infty}(M)^{G}$) gives rise to a closed embedding

### v2

@@ 5. Computing $\operatorname{Spec}(\mathscr{A})$
5.i. $\operatorname{Spec}(\mathscr{A})$ for a finitely generated $C^{\infty}$-ring $\mathscr{A}$
@@
Definition 5.1.
A $C^{\infty}$-ring $\mathscr{A}$ is finitely generated if there is a surjective map $\varpi:C^{\infty}(\mathbb{R}^{n})\to\mathscr{A}$ of $C^{\infty}$-rings. And then $\{\varpi(x_{1}),\ldots\varpi(x_{n})\}\subset\mathscr{A}$ is a set of generators of $\mathscr{A}$.
@@
5.ii. $\operatorname{Spec}(\mathscr{A})$ for a point determined $C^{\infty}$-ring $\mathscr{A}$
@@ A $C^{\infty}$-ring $\mathscr{A}$ is point determined if for every element $a\in\mathscr{A}$
$\big(x(a)=0\textrm{ for all }x\in\operatorname{Hom}(\mathscr{A},\mathbb{R})\equiv X_{\mathscr{A}}\big)\quad\Rightarrow a=0.$
@@
(5.2) $\mathscr{A}\to C^{0}(X_{\mathscr{A}}),\quad a\mapsto a_{*}$
@@
Suppose $a\in\bigcap_{x\in U}I_{x}$. Then for any $x\in U$, we have $a\in I_{x}$. This means that there is $d\in\mathscr{A}$ so that $x(d)\not=0$ and $ad=0$. Since $0\not=x(d)=d_{*}(x)$ and $d_{*}$ is continuous there is a neighborhood $V$ of $x$ so that $d_{*}(y)\not=0$ for all $y\in V$. Since $ad=0$, $0=(ad)_{*}=a_{*}d_{*}$. Hence $a_{*}|_{V}=0$. Since $x\in U$ is arbitrary, $a_{*}|_{U}=0$. Therefore
@@
$0=a_{*}d_{*}=(ad)_{*}$
Since the map (5.2) is injective ($\mathscr{A}$ is point determned), it follows that $da=0$ Since $x(d)=x_{*}(d)\not=0$ it follows that $a\in I_{x}$. Since $x\in U$ is arbitrary, $a\in\bigcap_{x\in U}I_{x}$. Therefore $\mathfrak{s}_{a}|_{U}=0$ and $\psi_{U}$ is injective. ∎
@@
$I(Z_{J})=\{f\in C^{\infty}(\mathbb{R}^{n})\mid f(p)=0\textrm{ for all }p\in Z_{J}\}$
@@
(5.4) $\overline{\operatorname{\mathsf{ev}}}:Z_{J}\to\operatorname{Hom}(C^{\infty}(\mathbb{R}^{n}/J),\mathbb{R}),\quad\overline{\operatorname{\mathsf{ev}}}_{p}(f+J)=f(p).$
@@
for all $p\in Z_{J}$. Hence $x(f+J)=0$ for all $x\in\operatorname{Hom}(\mathscr{A},\mathbb{R})$. Since $\mathscr{A}$ is point determined, $f+J=0$ in $\mathscr{A}=C^{\infty}(\mathbb{R}^{n})/J$. Thus $f\in J$. ∎
It follows from 5.i that if $\mathscr{A}=C^{\infty}(\mathbb{R}^{n})/J$ is point determined then $X_{\mathscr{A}}=Z_{J}$ and $\mathscr{O}_{A}$ is given by
@@
5.iii. The spectrum of a $C^{\infty}$-ring $\operatorname{Spec}(C^{\infty}(M)^{G})$ of invariant functions.
@@
If $f\in C^{\infty}(M)^{G}$ and $x(f)=0$ for all $x\in\operatorname{Hom}(C^{\infty}(M)^{G},\mathbb{R})$, then $0=\operatorname{\mathsf{ev}}_{p}(f)=f(p)$ for all $p\in M$. Therefore $f=0$. ∎
@@
(5.5) $\overline{\operatorname{\mathsf{ev}}}:M/G\to X_{C^{\infty}(M)^{G}},\quad\overline{\operatorname{\mathsf{ev}}}(G\cdot p)=\operatorname{\mathsf{ev}}_{p}.$
In general there is no reason for the map $\overline{\operatorname{\mathsf{ev}}}$ to be a bijection. If invariant functions separate orbits (which happens when the action is proper) then $\overline{\operatorname{\mathsf{ev}}}$ is injective. In general it is not. For instance consider the action of the reals $\mathbb{R}$ on the torus $\mathbb{T}^{2}$ by irrational flow (cf. Subsection 4.i). Then the ring of invariants $C^{\infty}(\mathbb{T}^{2})^{\mathbb{R}}$ consists of constant functions. Consequently it has exactly one $\mathbb{R}$ point and the map $\overline{\operatorname{\mathsf{ev}}}:\mathbb{T}^{2}/\mathbb{R}\to X_{C^{\infty}(\mathbb{T}^{2})^{\mathbb{R}}}$ is not injective. We now argue that for proper actions the map $\overline{\operatorname{\mathsf{ev}}}$ of (5.5) is a homeomorphism. We do it in a series of lemmas.
Lemma 5.7.
@@
Remark 5.8.
Let $G\times M\to M$ be an action of a Lie group $G$ on a manifold $M$. Let $q:M\to M/G$ denote the orbit/quotient map (in the category of topological spaces). By the universal property of $q$ for any $G$-invariant function $f:M\to\mathbb{R}$ there exists a unique continuous function $\overline{f}:M/G\to\mathbb{R}$ with $f=\overline{f}\circ q$. We thus get an injective map
(5.6) $\Upsilon:C^{\infty}(M)^{G}\to C^{0}(M/G),\qquad f\mapsto\overline{f}$
@@
(5.7) $C^{\infty}(M/G):=\Upsilon(C^{\infty}(M)^{G})\qquad$
@@
Lemma 5.9.
Let $G\times M\to M$ be a proper action of a Lie group $G$ on a manifold $M$. Then the $C^{\infty}$-subring $C^{\infty}(M/G)$ of $C^{0}(M/G)$ defined in Remark 5.8 contains a proper function. Consequently the map
@@
In the course of the proof of Theorem 4.3.1 in [Pa] Palais constructs a sequence of invariant functions $\{f_{n}\}_{n\in\mathbb{N}}\subset C^{\infty}(M)^{G}$ with the property that their supports form a locally finite cover of $M$, that $f_{n}(M)\subset[0,1]$ for all $n$, $\sum f_{n}=1$ and that sets $q(\operatorname{supp}(f_{n}))\subset M/G$ are compact (here as before $q:M\to M/G$ is the quotient map). Let $\overline{f}_{n}$ denote the image of $f_{n}$ in $C^{0}(M/G)$ under the map $\Upsilon:C^{\infty}(M)^{G}\to C^{0}(M/G)$ (cf. Remark 5.8). Then the function
$h=\sum_{n=1}^{\infty}n\overline{f}_{n}$
is a desired proper function (cf. proof of Proposition 2.28 in [Lee]).
Hence for any $\varphi\in X_{C^{\infty}(M/G)}\simeq X_{C^{\infty}(M)^{G}}$ there is an orbit $G\cdot m\in M/G$ so that $\varphi(\overline{f})=\overline{f}(G\cdot m)=f(m)$ for all $f\in C^{\infty}(M)^{G}$. Here as before $\overline{f}=\Upsilon(f)$. Hence the map $\overline{\operatorname{\mathsf{ev}}}$ is onto. We have already observed that since the action of $G$ on $M$ is proper, the map $\overline{\operatorname{\mathsf{ev}}}$ is injective. Thus $\overline{\operatorname{\mathsf{ev}}}$ is a bijection. ∎
Lemma 5.10.
@@
As we observed in Remark 5.8 the map $\Upsilon:C^{\infty}(M)^{G}\to C^{\infty}(M/G)$ is an isomorphism of $C^{\infty}$-rings, hence $X_{C^{\infty}(M)^{G}}$ and $X_{C^{\infty}(M/G}$ are homeomorphic. By Lemma 5.9 the evaluation map
@@
is a bijection. We need to show that $\operatorname{\mathsf{ev}}$ is a homeomorphism. By Lemma 5.11 below, the topology on $M/G$ is generated by the sets of the form $\{h\not=0\}:=\{p\in M/G\mid h(p)\not=0\}$, $h\in C^{\infty}(M/G)$. On the other hand the topology on $X_{C^{\infty}(M/G)}$ is generated by the sets of the form
@@
Lemma 5.11.
Let $G\times M\to M$ be a proper action of a Lie group $G$ on a manifold $M$, $q:M\to M/G$ the quotient map and $C^{\infty}(M/G)\subset C^{0}(M/G)$ defined by (5.7). The quotient topology on $M/G$ is the smallest topology making all the functions in $C^{\infty}(M/G)$ continuous. That is, the quotient topology is generated by the sets of the form $\{h\not=0\}$ for all $h\in C^{\infty}(M/G)$.
@@
To prove the existence of $f$ choose a point $m\in q^{-1}(p)\subset M$. Let $K$ denote the stabilizer of $m$. Since the action of $G$ is proper there is a slice $\Sigma$ through $m$ for the action. Then $G\cdot\Sigma$ is open in $M$, is diffeomorphic to $(G\times\Sigma)/K$ and $(G\cdot\Sigma)/G\simeq\Sigma/K$ as topological spaces. Consequently $q(G\cdot\Sigma)$ is an open neighborhood of $p$ in $M/G$ and the composite $\Sigma\hookrightarrow M\xrightarrow{q}M/G$ descends to an open embedding $q^{\prime}:\Sigma/K\to M/G$. On $\Sigma$ choose a smooth $K$-invariant function $\rho$ so that $\rho$ is identically 1 near $m$ and $\operatorname{supp}(\rho)$ is contained in a compact set which is disjoint from $(q^{\prime})^{-1}(C)$. Since $\rho$ is $K$-invariant, it descends to a continuous function $\tau$ on $\Sigma/K\simeq(G\cdot\Sigma)/G\hookrightarrow M/G$. Extend $\tau$ by 0 to all of $M/G$. The resulting function is the desired function $f$. ∎
Lemma 5.12.
@@
By Lemma 5.10 the space of points $X_{C^{\infty}(M)^{G}}$ is homeomorphic to the orbit space $M/G$ by way of the map
@@
(5.8) $(\overline{\operatorname{\mathsf{ev}}})^{*}:C^{0}(X_{C^{\infty}(M)^{G}})\to C^{0}(M/G),\quad h\mapsto h\circ\overline{\operatorname{\mathsf{ev}}}$
@@
By Lemma 5.6 the $C^{\infty}$-ring $C^{\infty}(M)^{G}$ is point determined, hence can be identified with a $C^{\infty}$-subring of $C^{0}(X_{C^{\infty}(M)^{G}})$ by way of
(5.9) $(\,\,)_{*}:C^{\infty}(M)^{G}\to C^{0}(X_{C^{\infty}(M)^{G}}),\quad f\mapsto f_{*}$
@@
the composite of (5.9) and (5.8) is the map $\Upsilon:C^{\infty}(M)^{G}\to C^{0}(M/G)$ that sends the invariant function $f\in C^{\infty}(M)^{G}$ to the unique function $\overline{f}:M/G\to\mathbb{R}$ with $\overline{f}\circ q=f$ (cf. Lemma 5.8). Lemma 5.3 now implies the desired result. ∎
5.iv. $\operatorname{Spec}(C^{\infty}(M)^{G})$ when the action $G\times M\to M$ is proper and $C^{\infty}(M)^{G}$ is finitely generated.
We now put together Subsections 5.i and 5.iii. Suppose now $G\times M\to M$ is a proper action and the $C^{\infty}$-ring of invariants $C^{\infty}(M)^{G}$ is finitely generated (hence there is a surjective map $\varpi:C^{\infty}(\mathbb{R}^{N})\to C^{\infty}(M)^{G}$ for some $N$). Then, one one hand,
@@
6. A remark on the Cushman-Weinstein conjecture
Suppose $M$ is a symplectic manifold with a proper Hamiltonian action of a Lie group $G$ and an associated equivariant moment map $\mu:M\to\mathfrak{g}^{\vee}$. Suppose further that the $C^{\infty}$-ring $C^{\infty}(M)^{G}$ is finitely generated. Then the $C^{\infty}$-ring of invariants $C^{\infty}(M)^{G}$ is Poisson and consequently $\operatorname{Spec}(C^{\infty}(M)^{G})\simeq(M/G,C^{\infty}_{M/G})$ is a Poisson $C^{\infty}$-scheme. Moreover as we saw in 5.iv a choice of a surjective map $\varpi:C^{\infty}(\mathbb{R}^{N})\to C^{\infty}(M)^{G}$ (i.e., a choice of generators of $C^{\infty}(M)^{G}$) gives rise to a closed embedding